# Med-R1: Reinforcement Learning for Generalizable Medical Reasoning in Vision-Language Models

### v4

@@ 4.1 Setup
Datasets. We adopt the VQA data from the open-access part of the OmniMedVQA benchmark [33], which consists of a total of $82,059$ images and $88,996$ vision question answering pairs. OmniMedVQA includes VQA pairs from eight imaging modalities: CT ($15,808$), MRI ($31,877$), X-Ray ($7,916$), Ultrasound ($10,991$), Dermoscopy ($6,679$), Fundus ($5,398$), OCT ($,64$ 646), and Microscopy ($5,680$). It is also categorized into five VQA question types, including Anatomy Identification ($16,448$), Disease Diagnosis ($55,387$), Lesion Grading ($2,098$), Modality Recognition ($11,565$), and Other Biological Attributes ($3,498$). We split the dataset into training and test sets following an 80-20 ratio for each setting.
@@
Baseline Methods & Evaluation Metric. Our evaluation framework integrates three classes of baselines: general VLMs including BLIP-2 [16], InstructBLIP [27], and the Qwen2-VL series (2B/7B/72B) [3, 30]; medical VLMs pretrained on domain-specific corpora (Med-Flamingo [6], MedVInT [7], LLaVA-Med [8]); and SFT variants of Qwen2-VL-2B.
We evaluate model performance using VQA choice accuracy, the standard metric for medical VQA, where models select answers from clinically validated options. Given a medical image $I$, a question $Q$, and $K$ candidate answers $\{A_{k}\}_{k=1}^{K}$ (one correct and $K-1$ distractors), accuracy is computed as:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\left(\hat{y}_{i}=y_{i}\right)$ (2)
where $N$ is the total number of test cases, $\hat{y}_{i}$ is the model’s predicted answer index, $y_{i}$ is the ground truth index, and $\mathbb{I}$ is the indicator function, which returns 1 if the predicted answer matches the ground truth and 0 otherwise.
@@ 4.2 Cross-Modality Generalization
Table 4: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition. The base model is Qwen2-VL-2B.
@@
Anatomy Identification 96.06 59.07 55.51 98.62 74.86 76.83
Disease Diagnosis 54.16 98.25 73.85 97.87 84.09 81.64
Lesion Grading 52.30 57.72 86.24 97.32 74.29 73.57
Modality Recognition 55.30 55.43 56.88 99.46 70.88 67.59
Other Attributes 56.60 59.93 56.88 97.91 96.59 73.58
Overall 62.88 66.08 65.87 98.24 80.14 74.64
Table 5: No-Think Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. The base model is Qwen2-VL-2B.
Train Test Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Anatomy Identification 96.33 59.29 64.22 99.04 76.56 79.09
Disease Diagnosis 55.00 98.30 63.76 98.08 85.23 80.07
Lesion Grading 54.22 58.04 96.33 97.37 72.30 75.65
Modality Recognition 56.84 57.95 55.05 99.79 73.15 68.56
Other Attributes 56.29 58.00 52.29 97.90 99.29 72.75
Overall 63.74 66.32 66.33 98.44 81.31 75.22
@@
[…]
@@
Qwen2-VL-2B (GRPO + Think) 62.88 66.08 65.87 98.24 80.14 74.64
Qwen2-VL-2B (GRPO + Nothink) 63.74 66.32 66.33 98.44 81.31 75.22
Qwen2.5-VL-3B (GRPO + Nothink) 55.47 68.72 80.00 96.87 85.51 77.31
Comparison to zero-shot and SFT evaluations with other VLMs. As demonstrated in Table 3, Med-R1 demonstrates its superiority across all eight medical imaging modalities while maintaining exceptional parameter efficiency. For zero-shot results, each cell denotes the zero-shot evaluation accuracy of the model on the particular modality. For the fine-tuned VLM results (last two rows), each cell represents the overall accuracy, reflecting the average generalization performance of the given modality when evaluated using models that were separately fine-tuned on each of the eight training modalities. Against general-purpose VLMs, our 2B-parameter model achieves 69.91% overall accuracy, surpassing the 72B-parameter Qwen2-VL by 1.86%—a notable result given the 36×36$\times$36 × parameter disparity. This advantage amplifies in critical diagnostic tasks: Med-R1 attains 71.67% accuracy in MRI compared to Qwen2-VL-72B’s 69.39%, and achieves 72.33% versus 65.31% in dermoscopy, challenging the prevailing scale-equals-performance paradigm. The limitations of specialized medical VLMs become evident through Med-Flamingo’s 30.38% average accuracy, which Med-R1 outperforms by 39.53%. This stark contrast underscores the ineffectiveness of narrow medical pretraining compared to our RL-driven adaptation strategy. When compared to supervised fine-tuning approaches, GRPO delivers 15.84% accuracy gains over SFT-tuned Qwen2-VL-2B (69.91% vs. 54.07%), with particularly significant improvements in CT interpretation (66.30% vs. 51.74%) and OCT analysis (71.96% vs. 53.99%).
@@ 4.3 Cross-Task Generalization
Figure 5: Rewards & Length. The top row shows the change in RL rewards during Med-R1 training across eight imaging modalities and five task types. We observe that the model quickly converges to stable reward values within 100–200 steps, indicating the computational efficiency of our RL post-training process. The bottom row shows the output completion length over training steps. Interestingly, the output length gradually decreases as training progresses, suggesting that the reward function discourages overly verbose or unfaithful reasoning. This observation challenges the conventional assumption that longer reasoning necessarily benefits vision-language performance in VQA.
4.4 Analysis of Thinking Process
@@
4.5 Analysis of Rewards and Completion Length.
The top row of Figure 5 shows the reward dynamics during GRPO-based reinforcement learning. Across all modalities and tasks, Med-R1 converges within 100–200 steps, as evidenced by the log-scale reward curves. This rapid and consistent convergence highlights the efficiency of our training pipeline and the model’s ability to quickly adapt to the reward structure. Moreover, the smooth reward trajectories suggest that GRPO remains stable even with limited RL updates, making it well-suited for low-resource scenarios.
Unlike in scientific domains such as mathematics [12], where model responses become increasingly detailed as thinking deepens, we observe a different trend in Med-RL training: the average response length decreases rapidly and then fluctuates (Figure 5). Rather than generating elaborate rationales, the model learns to produce shorter, more direct outputs—potentially favoring conciseness over verbosity. This aligns with our findings in subsection 4.4, suggesting that extended reasoning may not always benefit cross-domain VQA, especially under large domain shifts. We hypothesize that RL without CoT supervision encourages the model to rely on its pre-trained knowledge when generating rationales. While this may be effective in general domains, it can lead to hallucinated or misaligned reasoning in specialized domains like medical images. These results challenge the common assumption that more thinking always improves performance, and highlight that the “quality”—rather than the “presence”—of thinking is a crucial factor affecting the outcome of RL post-training.
@@
This work marks an initial step in applying RL to medical vision-language models. We adopt a frame-level VQA setting for consistent evaluation across modalities, but this simplifies the real-world complexity of medical imaging. In practice, CT and MRI are volumetric, and ultrasound is dynamic, requiring reasoning across slices and time. Future directions include extending Med-R1 to support multiple inputs. We also plan to incorporate patient context and explore more advanced reasoning frameworks for clinical deployment.
@@ 5 Conclusion
We present Med-R1, a RL-enhanced vision-language model that improves medical reasoning across diverse imaging modalities. By leveraging GRPO, Med-R1 surpasses the limitations of supervised fine-tuning, achieving strong cross-modality and cross-task generalization. Despite its compact 2B size, it matches or exceeds much larger models—including domain-specific and general-purpose VLMs—while remaining lightweight and deployable in resource-constrained clinical settings. Interestingly, we find that RL-based learning without explicit step-by-step rationales can outperform reasoning-heavy approaches, particularly under domain shift. This suggests that the quality, not quantity, of reasoning is key for robust medical adaptation. Med-R1 opens promising directions for developing scalable, reliable, and efficient AI systems in healthcare.

### v5

@@ 4.1 Setup
Datasets. We adopt the VQA data from the open-access part of the OmniMedVQA benchmark [36], which consists of a total of $82,059$ images and $88,996$ vision question answering pairs. OmniMedVQA includes VQA pairs from eight imaging modalities: CT ($15,808$), MRI ($31,877$), X-Ray ($7,916$), Ultrasound ($10,991$), Dermoscopy ($6,679$), Fundus ($5,398$), OCT ($4,646$), and Microscopy ($5,680$). It is also categorized into five VQA question types, including Anatomy Identification ($16,448$), Disease Diagnosis ($55,387$), Lesion Grading ($2,098$), Modality Recognition ($11,565$), and Other Biological Attributes ($3,498$). We split the dataset into training and test sets following an 80-20 ratio for each setting.
@@
Table 5: Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization. Results demonstrate that domain-specific training (e.g., Disease Diagnosis) preserves in-task expertise while maintaining adaptability to unseen tasks, particularly for modality-agnostic skills like Modality Recognition. The base model is Qwen2.5-VL-3B.
RL fine-tuned Qwen2.5-VL-3B
Train Test Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Anatomy Identification $94.65\pm 0.75$ $61.87\pm 0.90$ $63.30\pm 4.59$ $96.66\pm 0.73$ $78.69\pm 3.05$ $79.04\pm 0.65$
Disease Diagnosis $43.83\pm 1.67$ $97.74\pm 0.28$ $79.82\pm 3.67$ $96.66\pm 0.73$ $87.78\pm 2.41$ $81.16\pm 0.49$
Lesion Grading $47.10\pm 1.71$ $59.33\pm 0.93$ $88.99\pm 2.98$ $96.03\pm 0.77$ $69.03\pm 3.41$ $72.10\pm 0.70$
Modality Recognition $45.96\pm 1.67$ $58.63\pm 0.91$ $67.89\pm 4.36$ $99.08\pm 0.40$ $69.46\pm 3.41$ $68.20\pm 0.71$
Other Attributes $44.04\pm 1.68$ $60.09\pm 0.92$ $60.09\pm 4.59$ $95.74\pm 0.82$ $94.74\pm 1.63$ $70.94\pm 0.70$
Overall $55.12\pm 0.75$ $67.53\pm 0.39$ $72.02\pm 1.86$ $96.83\pm 0.31$ $79.94\pm 1.32$ $74.29\pm 0.65$
Table 6: No-Think Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
No-Think RL fine-tuned Qwen2.5-VL-3B
Train Test Anatomy Disease Lesion Modality Other Overall
Identification Diagnosis Grading Recognition Attributes
Anatomy Identification $96.85\pm 0.57$ $63.86\pm 0.90$ $79.59\pm 3.78$ $96.82\pm 0.69$ $77.41\pm 3.12$ $82.91\pm 0.64$
Disease Diagnosis $44.82\pm 1.73$ $98.78\pm 0.21$ $85.78\pm 3.33$ $95.82\pm 0.82$ $90.91\pm 2.13$ $83.22\pm 0.48$
Lesion Grading $44.61\pm 1.70$ $60.27\pm 0.91$ $97.02\pm 1.61$ $95.90\pm 0.79$ $82.10\pm 2.84$ $75.98\pm 0.70$
Modality Recognition $47.07\pm 1.71$ $59.58\pm 0.91$ $75.92\pm 3.90$ $99.75\pm 0.19$ $81.53\pm 2.84$ $72.77\pm 0.70$
Other Attributes $43.98\pm 1.71$ $61.11\pm 0.90$ $61.70\pm 4.59$ $96.07\pm 0.77$ $95.60\pm 1.49$ $71.69\pm 0.71$
Overall $55.46\pm 0.77$ $68.72\pm 0.39$ $80.00\pm 1.67$ $96.87\pm 0.31$ $85.51\pm 1.16$ $77.31\pm 0.62$
We focus on VQA, which integrates core vision–language abilities such as classification, grounding, and reasoning into a unified framework, making it a representative precursor to detection or captioning tasks.
Baseline Methods & Evaluation Metric.
We report our results by separating the baselines into three groups. Zero-shot VLMs are models pre-trained for general-purpose VQA without medical adaptation. Medical VLMs are models pre-trained on medical data specifically for medical VQA. Fine-tuned VLMs are models trained on OmniMedVQA using supervised fine-tuning and RL fine-tuning. We evaluate model performance using VQA choice accuracy, the standard metric for medical VQA, where the model selects the correct answer from $K$ clinically validated options. Given an image $I$, a question $Q$, and candidate answers $\{A_{k}\}_{k=1}^{K}$, accuracy is defined as:
$\text{Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(\hat{y}_{i}=y_{i}),$ (2)
where $N$ is the total number of test cases, $\hat{y}_{i}$ is the predicted answer index, $y_{i}$ the ground truth, and $\mathbb{I}$ the indicator function (1 if correct, 0 otherwise).
To assess result reliability, we computed 95% bootstrap confidence intervals (CIs) for all model variants across modalities and tasks. Each experiment performed 10,000 bootstrap resamplings of binary accuracy labels to estimate the mean and confidence bounds, reporting $mean\pm half\text{-}width$ for each model.
All fine-tuned VLM results (SFT, GRPO+Think, GRPO+Think-After, and GRPO+No-Think) are computed under the identical averaging protocol—macro-averaged accuracy across all cross-modality training–testing pairs—to ensure fair and consistent comparison.
@@ 4.2 Cross-Modality Generalization
Table 7: Think-After Cross-Task Generalization of Med-R1: Performance is evaluated across five clinical reasoning task types (rows: training tasks, columns: test tasks), with darker cell shading indicating stronger generalization.
Think-After fine-tuned Qwen2.5-VL-3B
@@
Anatomy Identification $95.22\pm 0.74$ $60.62\pm 0.90$ $65.83\pm 4.36$ $96.57\pm 0.73$ $81.53\pm 2.91$ $79.95\pm 0.65$
Disease Diagnosis $43.11\pm 1.68$ $97.86\pm 0.27$ $81.65\pm 3.67$ $95.99\pm 0.77$ $92.33\pm 1.99$ $82.19\pm 0.49$
Lesion Grading $43.56\pm 1.70$ $59.78\pm 0.92$ $85.55\pm 3.21$ $96.20\pm 0.77$ $78.55\pm 3.05$ $72.73\pm 0.71$
Modality Recognition $44.73\pm 1.71$ $58.82\pm 0.93$ $68.12\pm 4.47$ $98.87\pm 0.44$ $78.12\pm 3.05$ $69.73\pm 0.71$
Other Attributes $45.21\pm 1.67$ $59.65\pm 0.90$ $67.43\pm 4.36$ $95.74\pm 0.82$ $95.74\pm 1.49$ $72.75\pm 0.71$
Overall $54.36\pm 0.75$ $67.34\pm 0.40$ $73.72\pm 1.86$ $96.67\pm 0.32$ $85.26\pm 1.18$ $75.47\pm 0.63$
@@
$\text{LLaMA Adapter v2}^{\dagger}$ [29] 33.72 31.19 41.99 37.29 34.22 32.82
$\text{MiniGPT-4}^{\dagger}$ [30] 28.88 30.47 34.56 26.43 30.36 29.74
@@
Qwen2-VL-2B (Think) 62.88 66.08 65.87 98.24 80.14 74.64
Qwen2-VL-2B (Nothink) 63.74 66.32 66.33 98.44 81.31 75.22
Qwen2.5-VL-3B (Think) 55.12 67.53 72.02 96.83 79.94 74.29
Qwen2.5-VL-3B (Think-after) 54.36 67.34 73.72 96.67 85.26 75.47
Qwen2.5-VL-3B (Nothink) 55.47 68.72 80.00 96.87 85.51 77.31
Comparison to zero-shot and SFT evaluations with other VLMs. As demonstrated in Table 4, Med-R1 demonstrates its superiority across all eight medical imaging modalities while maintaining exceptional parameter efficiency. For zero-shot results, each cell denotes the zero-shot evaluation accuracy of the model on the particular modality. For all the fine-tuned VLM results, each cell represents the overall accuracy, reflecting the average generalization performance of the given modality when evaluated using models that were separately fine-tuned on each of the eight training modalities. Against general-purpose VLMs, our 2B-parameter model achieves 69.91% overall accuracy, surpassing the 72B-parameter Qwen2-VL by 1.86%—a notable result given the $36\times$ parameter disparity. This advantage amplifies in critical diagnostic tasks: Med-R1 attains 71.67% accuracy in MRI compared to Qwen2-VL-72B’s 69.39%, and achieves 72.33% versus 65.31% in dermoscopy, challenging the prevailing scale-equals-performance paradigm. The limitations of specialized medical VLMs become evident through Med-Flamingo’s 30.38% average accuracy, which Med-R1 outperforms by 39.53%. This stark contrast underscores the ineffectiveness of narrow medical pretraining compared to our RL-driven adaptation strategy. When compared to supervised fine-tuning approaches, GRPO delivers 15.84% accuracy gains over SFT-tuned Qwen2-VL-2B (69.91% vs. 54.07%), with particularly significant improvements in CT interpretation (66.30% vs. 51.74%) and OCT analysis (71.96% vs. 53.99%).
Figure 3: Training accuracy–reward curves of the Think, No-Think, and Think-After models on CT and MRI datasets.
@@ 4.3 Cross-Task Generalization
4.4 Analysis of No-Think
@@
4.5 Analysis of Think-after
We introduce the Think-After strategy primarily to address the practical need for achieving both high accuracy and interpretable reasoning in medical VLMs (Table 7, Table 3). Rather than serving solely as a control experiment, Think-After is designed to meet the dual objective of preserving predictive performance while providing reasoning traces that clinicians can review and validate.
Beyond satisfying the need for both accuracy and interpretability, Think-After also provides insight into why the No-Thinking strategy sometimes outperforms the conventional Thinking approach. As shown in Figure 3, Think-After achieves faster convergence and higher accuracy rewards than the Thinking variant, suggesting that generating reasoning tokens before the answer may disrupt the autoregressive generation process and hinder optimization. However, Think-After still performs slightly below No-Thinking, indicating that additional mechanisms—such as residual contextual coupling or reasoning-token noise—may also contribute to the performance gap. While Think-After does not completely resolve the question of why No-Thinking surpasses both pre- and post-answer reasoning, it fulfills the essential requirement of combining strong accuracy with interpretable outputs, and simultaneously offers valuable clues toward understanding the dynamics between reasoning generation and performance in medical VLMs.
Table 9: Reader Study Results. Three readers independently evaluated 100 correctly answered VQA samples for factual correctness (Quality, 1–5) and reasoning–answer consistency (Consistency, %).
Reader Quality Consistency (%)
A 4.23 93.0
B 4.19 92.0
C 4.34 95.0
Mean 4.25 93.3
  • Quality reflects factual and clinical correctness of model answers; Consistency measures logical alignment between reasoning and answers.
4.6 Reader Study
To further assess interpretability and clinical relevance, we conducted a reader study with three researchers experienced in medical imaging and vision–language models. Each independently evaluated 100 representative VQA samples generated by different model variants, covering diverse imaging modalities and reasoning types (Table 9). Evaluations focused on (1) the factual and clinical correctness of model answers and (2) the logical consistency between the reasoning (“Think”) and the predicted answer (“Answer”). Across readers, the Think-After model achieved the highest agreement and was consistently judged to produce reasoning that was both coherent and clinically sound, demonstrating improved interpretability and reasoning reliability for medical decision-support applications.
@@ 4.6 Reader Study
This work marks an initial step in applying RL to medical vision–language models. We adopt a frame-level VQA setting for consistent evaluation across modalities, but this simplifies the real-world complexity of medical imaging. In practice, CT and MRI are volumetric, and ultrasound is dynamic, requiring reasoning across slices and time. Another limitation lies in the reasoning supervision and data scale. Our findings suggest that the “No-Thinking” model occasionally outperforms reasoning-enabled variants, which may stem from the limited availability of high-quality medical reasoning data.
In domains where clinically faithful chain-of-thought (CoT) annotations are scarce, the model may not learn to benefit from explicit reasoning, leading the reasoning process to introduce noise rather than insight. We believe this reflects a broader limitation of current medical datasets rather than of reasoning itself. Future work should explore scaling up medically grounded CoT data and aligning RL rewards with clinically validated reasoning quality to fully realize the benefits of explicit reasoning. Future directions also include extending Med-R1 to support multi-frame or volumetric inputs, incorporating patient context, and investigating more advanced reasoning frameworks for clinical deployment.
@@ 5 Conclusion
We present Med-R1, a reinforcement learning–enhanced vision–language model for improving medical reasoning across diverse imaging modalities and clinical tasks. Leveraging GRPO-based post-training, Med-R1 achieves strong cross-modality and cross-task generalization, surpassing the limits of supervised fine-tuning. Despite its compact 3B scale, Med-R1 performs competitively with or better than larger medical and general VLMs, while remaining efficient for deployment.
Through systematic analysis, we find that removing explicit reasoning (No-Think) improves convergence and generalization, suggesting that reasoning quality, not quantity, drives medical performance. To balance interpretability and accuracy, we introduce the Think-After strategy, which decouples reasoning from answer generation and enhances clinical transparency without compromising accuracy. Overall, Med-R1 establishes a scalable framework for reinforcement learning in medical VLMs and offers new insights into how reasoning dynamics interact with domain generalization, paving the way toward reliable and interpretable medical AI systems.